If a shopkeeper marks the price of goods 50% more than their cost price and allows a discount of 40%, what is his gain% or loss%?
A)20% loss
B)20% gain
C)10% loss
D)10% gain

Answer by Theo(13342)

c = cost
p = price
d = discount rate (discount rate = discount percent divided by 100)
s = sale price
m = markup rate (markup rate = markup percent divided by 100)

your formula would be:

s = (1-d)*p where:

p = (1+m)*c

if you put these 2 formulas together, you get:

s = (1-d)*(1+m)*c

you have:
d = .40
m = .50

your formula becomes:
s = (.6)*(1.5)*c which becomes:
s = .9*c

.9 is equivalent to 90% of cost which represents a 10% loss.

let's see how that works in real life.

assume the cost is $100.00

you mark it up to make the price equal to $150.00

you give a 40% discount which brings the sale price down to $150.00 minus .40 * $150.00 which equals $150.00 minus 60 dollars which equals $90.00

your cost was $100.
your sale price was $90.00
you lost $10.00
that represents a loss of $10/$100 = 10%